OpenStudy (anonymous):

what is an equation of the line that passes through the point (-2,5) and is perpinduclar to the line whose equation is y= 1/2x +5 CAN SOMEONE HELP ME SOLVE IT AND TELL ME WHY?

OpenStudy (calculusfunctions):

If the new line is perpendicular to the given line, then the new line has a slope value that is the negative reciprocal of the slope value of the given line.

OpenStudy (anonymous):

The gradient of a line is the value of m when you look at the equation y=mx+c, so for this example it is 1/2. The perpendicular gradient is given as -1/m, so you can work out the perpendicular gradient is -2.

OpenStudy (calculusfunctions):

y = mx + b is the equation of any line that has slope m and y-intercept b. Your given line is y = (1/2)x + 5. Hence the slope value of the given line is 1/2. The negative reciprocal of 1/2 is -2. Thus a line perpendicular to the given line has a slope of -2. Hence we have y = -2x + b. Now we're told that this line must pass through the point (-2, 5). Simply substitute x = -2 and y = 5, and solve for b.

OpenStudy (anonymous):

(x,y) = (-2,5) y = -2x +p => 5 = -2(-2) +p => 5 = 4 + p => p = 1 => y = -2x +1

OpenStudy (anonymous):

Slope of a Perpinduclar Line = mP Slope of a Line = m mP = -1/m your Line : y = 1/2 x + 5 => slope of line (m=1/2) Slope of Perpinduclar Line to your Line : mP = -2

OpenStudy (anonymous):

did you get it its >> y = -2x +1 <<